# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 9.2 Absorption
Otherwise $f^{\prime}$ is a projection (again by the structure of the majority clone) and Proposition 9.4 implies that $h\rightsquigarrow_{B}\operatorname{proj}_{i}$ for some $i$. But then, again using Proposition 9.4, a $k$-shift of $h$ (i.e. a term obtained by cyclically shifting the arguments of $h$ by $k$ positions) is $\rightsquigarrow_{B}$ related to $\operatorname{proj}_{i+k\bmod p}$. Finally $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ by the definition of $s$ (and, again, Proposition 9.4). This finishes the proof of the theorem. ∎
@@
    for every $f\in\operatorname{Clo}(\mathbf{A})$ there exists $g\in\operatorname{Clo}(\{0,1\};\operatorname{maj})$ such that $f\rightsquigarrow_{B}g$.
@@
We will begin by showing that (c) is equivalent to (d) in every algebra. For the implication from (d) to (c) take any $n$-ary operation $f$ of algebra $\mathbf{A}$, and let $g$ be such that $f\rightsquigarrow_{B}g$. Take two tuples ${\mathbf{a}},{\mathbf{a}}^{\prime}$ such that $(a_{1},a_{1}^{\prime}),\dotsc,(a_{n},a_{n}^{\prime})\in R$. If $f({\mathbf{a}})\in B$ then $(f({\mathbf{a}}),f({\mathbf{a}}^{\prime}))\in R$ and we are done, so suppose $f({\mathbf{a}})\notin B$. Let ${\mathbf{x}}$ be the characteristic tuple of ${\mathbf{a}}$; by the definition of $\rightsquigarrow_{B}$ we have $g({\mathbf{x}})=0$. On the other hand, the definition of $R$ and the choice of ${\mathbf{a}},{\mathbf{a}}^{\prime}$ implies that ${\mathbf{x}}^{\prime}$ – the characteristic tuple of ${\mathbf{a}}^{\prime}$ – is greater than or equal to ${\mathbf{1}}-{\mathbf{x}}=(1-x_{1},\dotsc,1-x_{n})$. Since $g$ is in $\operatorname{Clo}(\{0,1\};\operatorname{maj})$, the clone of monotone selfdual operations, we get $g({\mathbf{x}}^{\prime})\geq g({\mathbf{1}}-{\mathbf{x}})={\mathbf{1}}-g({\mathbf{x}})=1$, therefore $f({\mathbf{a}}^{\prime})\in B$. This proves the first implication.
@@
For the implication from (c) to (d), note that to every $n$-ary operation $f\in\operatorname{Clo}(\mathbf{A})$ we can associate $g^{\prime}:\{0,1\}^{n}\rightarrow\{0,1\}$ by putting $g^{\prime}({\mathbf{x}})=1$ if and only if $f({\mathbf{a}})\in B$ for all ${\mathbf{a}}$ with characteristic tuple greater than or equal to ${\mathbf{x}}$. The operation $g^{\prime}$ is monotone by definition, and clearly $f\rightsquigarrow_{B}g^{\prime}$. From $g^{\prime}$ we define $g$ in the following way: for every ${\mathbf{x}}$ we put $g({\mathbf{x}})=g^{\prime}({\mathbf{x}})$ unless $g^{\prime}({\mathbf{x}})=g^{\prime}({\mathbf{1}}-{\mathbf{x}})=1$ and $x_{1}=0$ – in this case we put $g({\mathbf{x}})=0$. The operation $g$ is clearly monotone, and we obviously have $f\rightsquigarrow_{B}g$. We will show that $g$ is self-dual, which will conclude the proof of (d). Indeed, if $g$ is not self-dual then we have $g^{\prime}({\mathbf{x}})=g^{\prime}({\mathbf{1}}-{\mathbf{x}})=0$. From the definition of $g^{\prime}$ we have two tuples ${\mathbf{a}}$ and ${\mathbf{a}}^{\prime}$ with characteristic tuples greater than or equal to ${\mathbf{x}}$ and ${\mathbf{1}}-{\mathbf{x}}$ respectively and such that $f({\mathbf{a}}),f({\mathbf{a}}^{\prime})\in A\setminus B$. Note that for every $i$ we have $R(a_{i},a_{i}^{\prime})$, and by compatibility $(f({\mathbf{a}}),f({\mathbf{a}}^{\prime}))\in R$, but this contradicts the condition that both $f({\mathbf{a}})$ and $f({\mathbf{a}}^{\prime})$ are outside $B$.
@@
Next, we assume (d) and aim to prove (b). Again, no assumptions on the algebra are necessary. We set $\{0,1\}$ to be the universe of $\mathbf{C}$ and for every basic operation of $\mathbf{A}$, say $f^{\mathbf{A}}$, put $f^{\mathbf{C}}=g$ where $g$ is any operation such that $f\rightsquigarrow_{B}g$. Clearly $\operatorname{Clo}(\mathbf{C})\subseteq\operatorname{Clo}(\{0,1\};\operatorname{maj})$ and therefore $\mathbf{C}$ has no $2$-absorbing subuniverses. Moreover we claim that the relation $R=A\times\{0\}\cup B\times\{1\}$ is a subuniverse of $\mathbf{A}\times\mathbf{C}$. Take $(a_{1},y_{1}),\dotsc,(a_{n},y_{n})\in R$, any $n$-ary symbol $f$ and let ${\mathbf{x}}$ be the characteristic tuple of ${\mathbf{a}}$. If $f^{\mathbf{A}}({\mathbf{a}})\in B$ we are done, so suppose that $f^{\mathbf{A}}({\mathbf{a}})\in A\setminus B$ and therefore $f^{\mathbf{C}}({\mathbf{x}})=0$. Then ${\mathbf{y}}\leq{\mathbf{x}}$ and so $f^{\mathbf{C}}({\mathbf{y}})=0$, which completes the proof.
@@
For (b) implies (a) see Proposition 4.4 and for (a) implies (d) see Theorem 9.5. Finally (a) always implies (e). For the converse implication we will use Proposition 8.7. By assumption, $B=\{b\}\trianglelefteq\mathbf{A}$. Take any $a\in A\setminus B$ and consider $\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$. If $(a,a)$ is an element of this algebra, then there is a term operation $t$ satisfying $t(a,b)=a=t(b,a)$ and, by Proposition 5.3, the set $\{a,b\}$ is a subuniverse. Since $t$ acts on $\{a,b\}$ as a semilattice with absorbing element $a$, the subalgebra of $\mathbf{A}$ with universe $\{a,b\}$ (which is minimal Taylor by Proposition 5.4) is term-equivalent to the semilattice with absorbing element $a$, but then $\{b\}$ can not be an absorbing subuniverse. This contradiction shows that the assumption of Proposition 8.7 is satisfied and that $\{b\}$ $3$-absorbs $\mathbf{A}$ as required. ∎
@@
Proof.
The proof of the previous theorem shows that all of the items other than “$B$ is a Taylor center of $\mathbf{A}$” are equivalent to the witness of centrality in item (d). If $\operatorname{Clo}(\mathbf{C})=\operatorname{Clo}(\{0,1\},\operatorname{maj})$, then $B$ is a Taylor center. In the other case, where $\operatorname{Clo}(\mathbf{C})$ is the clone of projections, all of the term operations are $\rightsquigarrow_{B}$-related to a projection, so $B$ is projective and then strongly projective by Theorem 5.7. We define an algebra $\mathbf{D}$ in the signature of $\mathbf{A}$ term-equivalent to the rock-paper-scissors algebra by interpreting an $n$-ary symbol $f$ in $\mathbf{D}$ using an arbitrarily chosen term involving all the essential variables of $f^{\mathbf{A}}$, e.g., $f^{\mathbf{D}}(d_{1},\dots,d_{n})=\operatorname{winner}(d_{1},\operatorname{winner}(d_{2},\dots)\dots)$ if all the coordinates of $f^{\mathbf{A}}$ are essential. The witness for $B$ being a Taylor center can then be taken $R=(A\times\{\mbox{rock}\})\cup(B\times D)$. Indeed, $\mathbf{D}$ has no nontrivial 2-absorbing subuniverse and $R$ is a subuniverse of $\mathbf{A}\times\mathbf{D}$: if $f$ is an $n$-ary symbol and ${\mathbf{a}},{\mathbf{d}}$ tuples such that $(a_{i},d_{i})\in R$ for each $i$, then either $a_{i}\in B$ for some essential coordinate $i$ of $f^{\mathbf{A}}$ and then $f^{\mathbf{A}}({\mathbf{a}})\in B$ by strong projectivity, or $a_{i}\not\in B$ for all the essential coordinates $i$ of $f^{\mathbf{A}}$, but then $d_{i}=\operatorname{rock}$ for all essential coordinates, therefore $f^{\mathbf{D}}({\mathbf{d}})=\operatorname{rock}$ by idempotency (note that $f^{\mathbf{A}}$ and $f^{\mathbf{D}}$ have the same essential coordinates). In both cases $(f^{\mathbf{A}}({\mathbf{a}}),f^{\mathbf{D}}({\mathbf{d}}))\in R$. ∎
@@
Note that $t$ generates the whole clone of $\mathbf{A}$. For (2) and the first part of (1) we apply $t$ to a tuple consisting of elements from $B\cup C$. The result will be from $B$ if the majority of the arguments are from $B$ and similarly for $C$.
For the second part of (1) take $f(x,y,z)=t(x,\dotsc,x,y,\dotsc,y,z,\dotsc,z)$ where the number of $x$’s, $y$’s and $z$’s is roughly equal (more precisely, the number of $x$’s and $y$’s is at least $p/2$ and the same is true for the other two pairs). By construction $f$ is a witness both for $\ B\trianglelefteq_{3}\mathbf{A}$ and for $C\trianglelefteq_{3}\mathbf{A}$, and thus for $B\cap C$ as well.
To show (3), recall that absorption is transitive (see Subsection 7.5), so it is enough to verify that, for any $a\in A\setminus C$, the pair $(a,a)$ is not in $\operatorname{Sg}(C\times\{a\}\cup\{a\}\times C)$ and apply Proposition 8.7. If $a\not\in B$, this follows from the fact that the relation $B(x)\vee B(y)$ is a subuniverse of $\mathbf{A}^{2}$ (see Theorem 5.10) and if $a\in B\setminus C$, this follows from the fact that the relation $C(x)\vee C(y)$ on $B$ is a subuniverse of $\mathbf{B}^{2}$. ∎
@@ 9.3 Edges
Proof.
In (a) there is a binary term operation acting like the semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$. By Proposition 5.3 together with Proposition 5.4, the set $\{a/\theta,b/\theta\}$ is a subuniverse of $\mathbf{E}/\theta$ and thus is equal to it. By the classification of Post [55] and minimality of $\mathbf{E}/\theta$ we conclude (a). Case (b) is identical, except the operation is a ternary majority. In (c), Theorem 2.4 implies that we have the Mal’cev operation $x-y+z$ for some abelian group $\mathbf{G}$, so $\mathbf{E}/\theta$ is term equivalent to the affine Mal’cev algebra of $\mathbf{G}$. Since $E/\theta=\operatorname{Sg}_{\mathbf{E}/\theta}(a/\theta,b/\theta)=a/\theta+\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ (where $+$ and $\cdot$ are computed in $\mathbf{G}$), the set $\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ covers $G$, therefore the group homomorphism $\mathbb{Z}\to\mathbf{G}$ defined by $t\mapsto t(b/\theta-a/\theta)$ is surjective. We conclude using the first isomorphism theorem that $\mathbf{G}$ is isomorphic to $\mathbb{Z}/m$ for some $m$.
@@
It remains to obtain an operation acting as a semilattice on both semilattice edges. In the semilattice case both $b/\theta$ and $b/\gamma$ are 2-absorbing subuniverses of $\mathbf{E}$ and, by Theorem 5.7, e.g. $t(x,y,\dots y)$ for a cyclic $t$ is such an operation. In the majority case the reasoning is identical except that all the sets $a/\theta,b/\theta,a/\gamma,b/\gamma$ are 3-absorbing and all these absorptions can be witnessed by a single term (e.g. the term obtained by identifying variables in a cyclic term as in the proof of Proposition 5.12). ∎
@@
Proof.
Let $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$ and let $\theta$ be a maximal congruence on $\mathbf{E}$ witnessing the minimal edge. Since the only simple affine Mal’cev algebras are (up to isomorphism) the affine Mal’cev algebras of $\mathbb{Z}/{p}$ for a prime $p$, we have by Theorem 5.13 that $\mathbf{E}/\theta$ is term-equivalent to a two-element semilattice, a two-element majority algebra, or an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. None of these algebras have a nontrivial subalgebra and so, by minimality, there is no congruence incomparable with $\theta$. Indeed, if a congruence $\alpha$ is incomparable with the maximal congruence $\theta$, then at least one $\alpha$-class $B$ intersects two of the $\theta$-classes. Since $B/\theta\leq\mathbf{E}/\theta$ and $\mathbf{E}/\theta$ does not have nontrivial subalgebras, then $B$ intersects all of the classes; in particular the classes containing $a$ and $b$. By minimality of the edge $(a,b)$, we get $B=E$, a contradiction. Similarly, since $\theta\subseteq\mu_{E}$, also $\mu_{E}=\theta$. Indeed, otherwise by definition of $\mu_{E}$ there is a proper subuniverse $B$ of $\mathbf{E}$ intersecting two of the $\theta$-classes, which leads to a contradiction as above. ∎
@@
Let $\mathbf{A}$ be minimal Taylor algebra:
@@
    If $\mathbf{B}\trianglelefteq_{2}\mathbf{A}$ and $a\in A\setminus B$ there exists $b\in B$ such that $(a,b)$ is a semilattice edge and $\operatorname{Sg}_{\mathbf{A}}(a,b)=\{a,b\}$.
@@
    If $\mathbf{B}\trianglelefteq_{3}\mathbf{A}$ and $a\in A\setminus B$ there exists $b\in B$ such that:
@@
      $(a,b)$ is a semilattice edge and $\operatorname{Sg}_{\mathbf{A}}(a,b)=\{a,b\}$, or
@@
      $(a,b)$ is a majority edge and $\operatorname{Sg}_{\mathbf{A}}(a,b)\cap B$ is one of the two equivalence classes of the congruence witnessing it.
Proof.
We will prove both items simultaneously. First we choose $b\in B$ such that $\operatorname{Sg}_{\mathbf{A}}(a,b)$ is minimal under inclusion in the set $\{\operatorname{Sg}_{\mathbf{A}}(a,b)\}_{b\in B}$. We let $\mathbf{C}=\operatorname{Sg}_{\mathbf{A}}(a,b)$ and $D=C\cap B$, and note that by our choice of $b$, for every $b^{\prime}\in D$ we have $C=\operatorname{Sg}_{\mathbf{A}}(a,b^{\prime})$.
We put $R=\operatorname{Sg}_{\mathbf{C}^{2}}((a,b),(b,a))$. The first step is to show that if $D^{2}\cap R\neq\emptyset$ then $(b,b)\in R$ and $(a,b)$ is semilattice edge such that $\operatorname{Sg}_{\mathbf{C}}(a,b)=\{a,b\}$. Indeed, if $D^{2}\cap R\neq\emptyset$ then $(D+R)\cap D\neq\emptyset$ and $a\in D+R$ and consequently $D+R=C$. Further, if $D+R=C$ then there exists $b^{\prime}\in D$ such that $(b^{\prime},b)\in R$, but then $b-R$ contains both $a$ and $b^{\prime}$ and thus $b-R=C$; in particular $(b,b)\in R$. We have shown that in this case there is a binary term operation acting on $\{a,b\}$ as a join-semilattice operation with top $b$. By Proposition 5.3, $\{a,b\}$ is a subuniverse of $\mathbf{A}$, by Proposition 5.4 the subalgebra with this subuniverse is a minimal Taylor algebra, which is clearly term equivalent to a two-element semilattice.
Going back to the claims in the proposition, if $B\trianglelefteq_{2}\mathbf{A}$, then a pair in $D^{2}\cap R$ is produced by a single application of the operation witnessing the 2-absorption to $(a,b)$ and $(b,a)$. This finishes the 2-absorption case.
We are left with the case of $B\trianglelefteq_{3}\mathbf{A}$ and $R\cap D^{2}=\emptyset$. Note that $D\trianglelefteq_{3}\mathbf{C}$ and, by Theorem 5.10, $R\subseteq B(x)\vee B(y)$. Therefore $D+R=C\setminus D$ is a 3-absorbing subuniverse of $\mathbf{C}$. By item (2) in Proposition 5.12, we conclude that the partition of $C$ into $D$ and $C\setminus D$ defines a congruence of $\mathbf{C}$ and the quotient modulo this congruence is term equivalent to a two-element majority algebra. This finishes the proof. ∎
@@
Proof.
Let $(a,b)$ be as in the statement. Put $\mathbf{A}^{\prime}=\operatorname{Sg}_{\mathbf{A}}(a,b)$ and $\theta$ a congruence witnessing the edge. The subalgebra $b/\theta$ 2-absorbs $\mathbf{A}^{\prime}$, and Proposition 9.7 provides $b^{\prime}$ in $b/\theta$ such that $\operatorname{Sg}_{\mathbf{A}^{\prime}}(a,b^{\prime})=\{a,b^{\prime}\}$. Since $(a,b)$ is a minimal edge, we get $b=b^{\prime}$. ∎

### v6

@@ 9.2 Absorption
Otherwise $f^{\prime}$ is a projection (again by the structure of the majority clone) and Proposition 9.12 implies that $h\rightsquigarrow_{B}\operatorname{proj}_{i}$ for some $i$. But then, again using Proposition 9.12, a $k$-shift of $h$ (i.e. a term obtained by cyclically shifting the arguments of $h$ by $k$ positions) is $\rightsquigarrow_{B}$ related to $\operatorname{proj}_{i+k\bmod p}$. Finally $s\rightsquigarrow_{B}\operatorname{maj}_{p}$ by the definition of $s$ (and, again, Proposition 9.12). This finishes the proof of the theorem.
@@
    for every $f\in\operatorname{Clo}($ there exists $g\in\operatorname{Clo}(\{0,1\};\operatorname{maj})$ such that $f\rightsquigarrow_{B}g$.
@@
We will begin by showing that (c) is equivalent to (d) in every algebra. For the implication from (d) to (c) take any $n$-ary operation $f$ of algebra , and let $g$ be such that $f\rightsquigarrow_{B}g$. Take two tuples ${\mathbf{a}},{\mathbf{a}}^{\prime}$ such that $(a_{1},a_{1}^{\prime}),\dotsc,(a_{n},a_{n}^{\prime})\in R$. If $f({\mathbf{a}})\in B$ then $(f({\mathbf{a}}),f({\mathbf{a}}^{\prime}))\in R$ and we are done, so suppose $f({\mathbf{a}})\notin B$. Let ${\mathbf{x}}$ be the characteristic tuple of ${\mathbf{a}}$; by the definition of $\rightsquigarrow_{B}$ we have $g({\mathbf{x}})=0$. On the other hand, the definition of $R$ and the choice of ${\mathbf{a}},{\mathbf{a}}^{\prime}$ implies that ${\mathbf{x}}^{\prime}$ – the characteristic tuple of ${\mathbf{a}}^{\prime}$ – is greater than or equal to ${\mathbf{1}}-{\mathbf{x}}=(1-x_{1},\dotsc,1-x_{n})$. Since $g$ is in $\operatorname{Clo}(\{0,1\};\operatorname{maj})$, the clone of monotone selfdual operations, we get $g({\mathbf{x}}^{\prime})\geq g({\mathbf{1}}-{\mathbf{x}})={\mathbf{1}}-g({\mathbf{x}})=1$, therefore $f({\mathbf{a}}^{\prime})\in B$. This proves the first implication.
@@
For the implication from (c) to (d), note that to every $n$-ary operation $f\in\operatorname{Clo}($ we can associate $g^{\prime}:\{0,1\}^{n}\rightarrow\{0,1\}$ by putting $g^{\prime}({\mathbf{x}})=1$ if and only if $f({\mathbf{a}})\in B$ for all ${\mathbf{a}}$ with characteristic tuple greater than or equal to ${\mathbf{x}}$. The operation $g^{\prime}$ is monotone by definition, and clearly $f\rightsquigarrow_{B}g^{\prime}$. From $g^{\prime}$ we define $g$ in the following way: for every ${\mathbf{x}}$ we put $g({\mathbf{x}})=g^{\prime}({\mathbf{x}})$ unless $g^{\prime}({\mathbf{x}})=g^{\prime}({\mathbf{1}}-{\mathbf{x}})=1$ and $x_{1}=0$ – in this case we put $g({\mathbf{x}})=0$. The operation $g$ is clearly monotone, and we obviously have $f\rightsquigarrow_{B}g$. We will show that $g$ is self-dual, which will conclude the proof of (d). Indeed, if $g$ is not self-dual then we have $g^{\prime}({\mathbf{x}})=g^{\prime}({\mathbf{1}}-{\mathbf{x}})=0$. From the definition of $g^{\prime}$ we have two tuples ${\mathbf{a}}$ and ${\mathbf{a}}^{\prime}$ with characteristic tuples greater than or equal to ${\mathbf{x}}$ and ${\mathbf{1}}-{\mathbf{x}}$ respectively and such that $f({\mathbf{a}}),f({\mathbf{a}}^{\prime})\in A\setminus B$. Note that for every $i$ we have $R(a_{i},a_{i}^{\prime})$, and by compatibility $(f({\mathbf{a}}),f({\mathbf{a}}^{\prime}))\in R$, but this contradicts the condition that both $f({\mathbf{a}})$ and $f({\mathbf{a}}^{\prime})$ are outside $B$.
@@
Next, we assume (d) and aim to prove (b). Again, no assumptions on the algebra are necessary. We set $\{0,1\}$ to be the universe of and for every basic operation of , say $f$, put $f=g$ where $g$ is any operation such that $f\rightsquigarrow_{B}g$. Clearly $\operatorname{Clo}($ and therefore has no $2$-absorbing subuniverses. Moreover, we claim that the relation $R=A\times\{0\}\cup B\times\{1\}$ is a subuniverse of . Take $(a_{1},y_{1}),\dotsc,(a_{n},y_{n})\in R$, any $n$-ary symbol $f$ and let ${\mathbf{x}}$ be the characteristic tuple of ${\mathbf{a}}$. If $f({\mathbf{a}})\in B$ we are done, so suppose that $f({\mathbf{a}})\in A\setminus B$ and therefore $f({\mathbf{x}})=0$. Then ${\mathbf{y}}\leq{\mathbf{x}}$ and so $f({\mathbf{y}})=0$, which completes the proof.
@@
For (b) implies (a) see Proposition 4.4 and for (a) implies (d) see Theorem 9.14. Finally (a) always implies (e). For the converse implication we will use Proposition 8.12. By assumption, $B=\{b\}\trianglelefteq$. Take any $a\in A\setminus B$ and consider $\operatorname{Sg}((a,b),(b,a))$. If $(a,a)$ is an element of this algebra, then there is a term operation $t$ satisfying $t(a,b)=a=t(b,a)$ and, by Proposition 5.3, the set $\{a,b\}$ is a subuniverse. Since $t$ acts on $\{a,b\}$ as a semilattice with absorbing element $a$, the subalgebra of with universe $\{a,b\}$ (which is minimal Taylor by Proposition 5.4) is term-equivalent to the semilattice with absorbing element $a$, but then $\{b\}$ can not be an absorbing subuniverse. This contradiction shows that the assumption of Proposition 8.12 is satisfied and that $\{b\}$ $3$-absorbs as required.
@@
Proof 9.18.
The proof of the previous theorem shows that all of the items other than “$B$ is a Taylor center of ” are equivalent to the witness of centrality in item (d). If $\operatorname{Clo}($, then $B$ is a Taylor center. In the other case, where $\operatorname{Clo}($ is the clone of projections, all of the term operations are $\rightsquigarrow_{B}$-related to a projection, so $B$ is projective and then strongly projective by Theorem 5.7. We define an algebra in the signature of term-equivalent to the rock-paper-scissors algebra by interpreting an $n$-ary symbol $f$ in using an arbitrarily chosen term involving all the essential variables of $f$, e.g., $f(d_{1},\dots,d_{n})=\operatorname{winner}(d_{1},\operatorname{winner}(d_{2},\dots)\dots)$ if all the coordinates of $f$ are essential. The witness for $B$ being a Taylor center can then be taken $R=(A\times\{\mbox{rock}\})\cup(B\times D)$. Indeed, has no nontrivial 2-absorbing subuniverse and $R$ is a subuniverse of : if $f$ is an $n$-ary symbol and ${\mathbf{a}},{\mathbf{d}}$ tuples such that $(a_{i},d_{i})\in R$ for each $i$, then either $a_{i}\in B$ for some essential coordinate $i$ of $f$ and then $f({\mathbf{a}})\in B$ by strong projectivity, or $a_{i}\not\in B$ for all the essential coordinates $i$ of $f$, but then $d_{i}=\operatorname{rock}$ for all essential coordinates, therefore $f({\mathbf{d}})=\operatorname{rock}$ by idempotency (note that $f$ and $f$ have the same essential coordinates). In both cases $(f({\mathbf{a}}),f({\mathbf{d}}))\in R$.
@@
Note that $t$ generates the whole clone of . For (2) and the first part of (1) we apply $t$ to a tuple consisting of elements from $B\cup C$. The result will be from $B$ if the majority of the arguments are from $B$ and similarly for $C$.
For the second part of (1) take $f(x,y,z)=t(x,\dotsc,x,y,\dotsc,y,z,\dotsc,z)$ where the number of $x$’s, $y$’s and $z$’s is roughly equal (more precisely, the number of $x$’s and $y$’s is at least $p/2$ and the same is true for the other two pairs). By construction $f$ is a witness both for $\ B\trianglelefteq_{3}$ and for $C\trianglelefteq_{3}$, and thus for $B\cap C$ as well.
To show (3), recall that absorption is transitive (see Subsection 7.5), so it is enough to verify that, for any $a\in A\setminus C$, the pair $(a,a)$ is not in $\operatorname{Sg}(C\times\{a\}\cup\{a\}\times C)$ and apply Proposition 8.12. If $a\not\in B$, this follows from the fact that the relation $B(x)\vee B(y)$ is a subuniverse of $\mathbf{A}^{2}$ (see Theorem 5.10) and if $a\in B\setminus C$, this follows from the fact that the relation $C(x)\vee C(y)$ on $B$ is a subuniverse of $\mathbf{B}^{2}$.
@@ 9.3 Edges
Proof 9.20.
In case (a) there is a binary term operation acting like the semilattice operation on $\{a/\theta,b/\theta\}$ with top element $b/\theta$. By Proposition 5.3 together with Proposition 5.4, the set $\{a/\theta,b/\theta\}$ is a subuniverse of and thus is equal to it. By the classification of Post [post41:clones] and minimality of we conclude (a). Case (b) is identical, except the operation is a ternary majority. In (c), Theorem 2.4 implies that we have the Mal’cev operation $x-y+z$ for some abelian group , so is term equivalent to the affine Mal’cev algebra of . Since $E/\theta=\operatorname{Sg}(a/\theta,b/\theta)=a/\theta+\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ (where $+$ and $\cdot$ are computed in ), the set $\{t\cdot(b/\theta-a/\theta):t\in\mathbb{Z}\}$ covers $G$, therefore the group homomorphism $\mathbb{Z}\to$ defined by $t\mapsto t(b/\theta-a/\theta)$ is surjective. We conclude using the first isomorphism theorem that is isomorphic to $\mathbb{Z}/m$ for some $m$.
@@
It remains to obtain an operation acting as a semilattice on both semilattice edges. In the semilattice case both $b/\theta$ and $b/\gamma$ are 2-absorbing subuniverses of and, by Theorem 5.7, e.g. $t(x,y,\dots y)$ for a cyclic $t$ is such an operation. In the majority case the reasoning is identical except that all the sets $a/\theta,b/\theta,a/\gamma,b/\gamma$ are 3-absorbing and all these absorptions can be witnessed by a single term (e.g. the term obtained by identifying variables in a cyclic term as in the proof of Proposition 5.12).
@@
Proof 9.21.
Let and let $\theta$ be a maximal congruence on witnessing the minimal edge. Since the only simple affine Mal’cev algebras are (up to isomorphism) the affine Mal’cev algebras of $\mathbb{Z}/{p}$ for a prime $p$, we have by Theorem 5.13 that is term-equivalent to a two-element semilattice, a two-element majority algebra, or an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. None of these algebras have a nontrivial subalgebra and so, by minimality, there is no congruence incomparable with $\theta$. Indeed, if a congruence $\alpha$ is incomparable with the maximal congruence $\theta$, then at least one $\alpha$-class $B$ intersects two of the $\theta$-classes. Since $B/\theta\leq$ and does not have nontrivial subalgebras, then $B$ intersects all of the classes; in particular the classes containing $a$ and $b$. By minimality of the edge $(a,b)$, we get $B=E$, a contradiction. Similarly, since $\theta\subseteq\mu_{E}$, also $\mu_{E}=\theta$. Indeed, otherwise by definition of $\mu_{E}$ there is a proper subuniverse $B$ of intersecting two of the $\theta$-classes, which leads to a contradiction as above.
@@
Let be minimal Taylor algebra:
@@
    If and $a\in A\setminus B$ there exists $b\in B$ such that $(a,b)$ is a semilattice edge and $\operatorname{Sg}(a,b)=\{a,b\}$.
@@
    If and $a\in A\setminus B$ there exists $b\in B$ such that:
@@
      $(a,b)$ is a semilattice edge and $\operatorname{Sg}(a,b)=\{a,b\}$, or
@@
      $(a,b)$ is a majority edge and $\operatorname{Sg}(a,b)\cap B$ is one of the two equivalence classes of the congruence witnessing it.
Proof 9.23.
We will prove both items simultaneously. First we choose $b\in B$ such that $\operatorname{Sg}(a,b)$ is minimal under inclusion in the set $\{\operatorname{Sg}(a,b)\}_{b\in B}$. We let and $D=C\cap B$, and note that by our choice of $b$, for every $b^{\prime}\in D$ we have $C=\operatorname{Sg}(a,b^{\prime})$.
We put $R=\operatorname{Sg}((a,b),(b,a))$. The first step is to show that if $D^{2}\cap R\neq\emptyset$ then $(b,b)\in R$ and $(a,b)$ is semilattice edge such that $\operatorname{Sg}(a,b)=\{a,b\}$. Indeed, if $D^{2}\cap R\neq\emptyset$ then $(D+R)\cap D\neq\emptyset$ and $a\in D+R$ and consequently $D+R=C$. Further, if $D+R=C$ then there exists $b^{\prime}\in D$ such that $(b^{\prime},b)\in R$, but then $b-R$ contains both $a$ and $b^{\prime}$ and thus $b-R=C$; in particular $(b,b)\in R$. We have shown that in this case there is a binary term operation acting on $\{a,b\}$ as a join-semilattice operation with top $b$. By Proposition 5.3, $\{a,b\}$ is a subuniverse of , by Proposition 5.4 the subalgebra with this subuniverse is a minimal Taylor algebra, which is clearly term equivalent to a two-element semilattice.
Going back to the claims in the proposition, if $B\trianglelefteq_{2}$, then a pair in $D^{2}\cap R$ is produced by a single application of the operation witnessing the 2-absorption to $(a,b)$ and $(b,a)$. This finishes the 2-absorption case.
We are left with the case of $B\trianglelefteq_{3}$ and $R\cap D^{2}=\emptyset$. Note that $D\trianglelefteq_{3}$ and, by Theorem 5.10, $R\subseteq B(x)\vee B(y)$. Therefore $D+R=C\setminus D$ is a 3-absorbing subuniverse of . By item (2) in Proposition 5.12, we conclude that the partition of $C$ into $D$ and $C\setminus D$ defines a congruence of and the quotient modulo this congruence is term equivalent to a two-element majority algebra. This finishes the proof.
@@
Proof 9.24.
Let $(a,b)$ be as in the statement. Put and $\theta$ a congruence witnessing the edge. The subalgebra $b/\theta$ 2-absorbs , and Proposition 9.22 provides $b^{\prime}$ in $b/\theta$ such that $\operatorname{Sg}(a,b^{\prime})=\{a,b^{\prime}\}$. Since $(a,b)$ is a minimal edge, we get $b=b^{\prime}$.